The numerator of a fraction is 2 less than the denominator. if 3 is added both numerator and denominator, the fraction becomes 3/4 . Find the fraction
Answer by TimothyLamb(4379)

---
x/y
---
x = y - 2
(x + 3)/(y + 3) = 3/4
---
put the system of linear equations into standard form
---
x - y = -2
4(x + 3) = 3(y + 3)
---
x - y = -2
4x + 12 = 3y + 9
---
x - y = -2
4x - 3y = -3
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
x = 3
y = 5
---
system matrix
1 -1
4 -3
---
inverse of system matrix
-3 1
-4 1
---
determinant of system matrix = 1
---
answer:
the fraction = 3/5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php